OpenStudy (anonymous):

Using the formula,s = square root 24d , where s = speed and d = distance to stop in feet, find the speed of a car in miles per hour if it took 150 feet to stop. Round your answer to a whole number. A. 40 mph B. 60 mph C. 75 mph

Directrix (directrix):

s = sqrt (24d) s^2 = 24 * 150 s^2 = 3600 s = 60